### v1

# Language Models Can Predict Their Own Behavior

Abstract
Autoregressive Language Models output text by sequentially predicting the next token to generate, with modern methods like Chain-of-Thought (CoT) prompting achieving state-of-the-art reasoning capabilities by scaling the number of generated tokens. However, are there times when we can infer how the model will behave (e.g. abstain from answering a question) early in the computation, making generation unnecessary? We show that internal representation of input tokens alone can often precisely predict, not just the next token, but eventual behavior over the entire output sequence. We leverage this capacity and learn probes on internal states to create early warning (and exit) systems. Specifically, if the probes can confidently estimate the way the LM is going to behave, then the system will avoid generating tokens altogether and return the estimated behavior instead. On 27 text classification datasets spanning five different tasks, we apply this method to estimate the eventual answer of an LM under CoT prompting, reducing inference costs by 65% (average) while suffering an accuracy loss of no more than 1.4% (worst case). We demonstrate the potential of this method to pre-emptively identify when a model will abstain from answering a question, fail to follow output format specifications, or give a low-confidence response. We explore the limits of this capability, showing that probes generalize to unseen datasets, but perform worse when LM outputs are longer and struggle to predict properties that require access to knowledge that the models themselves lack. Encouragingly, performance scales with model size, suggesting applicability to the largest of models.  [^1] https://github.com/DhananjayAshok/LMBehaviorEstimation
1 Introduction
Language Models trained to predict the next token of a sentence on large-scale data [58, 14] are unrivalled in general Natural Language Understanding (NLU). During inference, these models generate remarkably high-quality outputs by sequentially predicting the most appropriate output tokens to generate. Frontier LM systems of today build on this by scaling inference i.e. increasing the number of tokens or sequences generated [64, 67, 48, 20], to provide even better responses.
In this paper we ask whether LMs have learned representations that contain information, not just about the next token the model will predict, but the entire sequence as a whole. We conclude with an affirmative answer, showing that the internal states of the LM can often be used to preemptively identify how an LM will behave. This information becomes accessible before the LM has generated a single token, potentially obviating the need for further inference on those instances.
Concretely (Figure 1), we train linear classifiers (probes) [3] that use the internal representation of the input tokens to predict the eventual behavior of the LM (e.g. whether an LM will comply with a request or abstain). We calibrate the probes using methods from conformal prediction [55], allowing them to estimate behavior only when confident. During inference, probes predict only when there is a provable guarantee on the estimation error, ensuring precise early warning signals for various model behaviors.
We apply this method to an LM that tackles text classification by generating several explanation tokens before a class prediction (CoT). The probes estimate the final CoT prediction and exit early if confident in their estimate. On 27 datasets, spanning the tasks of MCQA, Sentiment Analysis, Topic Classification, Toxicity Detection and Fact Verification, our method reduces inference costs by 65% on average, while suffering accuracy losses of no more than 1.4%. Furthermore, on 14/27 of the datasets, our method reduces inference costs by 63% on average with no cost to accuracy. The probes generalize to unseen datasets, reducing the inference cost of CoT on OpenBookQA [44] by 68% with minimal loss to accuracy (0.4%), despite only training on other MCQA datasets.
Beyond accelerating text classification, probes can pre-emptive identify several behaviors of interest. We apply our method to a QA system that verbalizes some reasoning about the input question and then either answers or abstains if the question is deemed to be unanswerable. The probes, which are trained to estimate whether the LM will abstain, make confident predictions on over 15% of the test instances and achieve an estimation consistency of over 90% when doing so.
Probes can provide precise early warning signals for degenerate behavior (whether the model will fail to follow a specified output format) and even ‘self-reflective" properties which cannot be inferred from the text output alone (whether the model will output a high perplexity answer to a question). While the number of instances where the probes give confident predictions varies significantly across datasets and tasks, the probe maintains high consistency on confident predictions.
We show that often fewer than a thousand training instances are sufficient for the probes to attain high estimation consistency, while ablations on the probing layer present a more complicated, task-specific story. Encouragingly, increasing the scale of the LM improves the performance of the probe, suggesting the method may scale favorably with ever-increasing model sizes [10, 24, 11, 16].
The probes are limited in the behaviors they can preemptively detect. While probes can identify various model behaviors, they fail on those that cannot be identified without knowledge external to the model (e.g. identifying whether the model is likely to output an incorrect answer to an MCQ instance). Additionally, probes are less confident on instances which trigger longer outputs.
We hope our findings enable practitioners to build efficient early warning systems on Language Models and enable further research into the information encoded in their hidden states [50, 7, 47].
Figure 1: Overview of our method: During training we collect the hidden states of the input token and eventual output behavior on a variety of prompts, using them to train a probe on the hidden states. With a validation set, we calibrate the probes for conformal prediction by identifying a threshold for confident predictions. During inference, if the probes are confident in their estimate of the eventual LM behavior, we stop all computation and return the prediction.
2 Background
Hidden State Probing: Lightweight probes have long been used to interpret the internal activations of neural networks [3] and language models [50, 7]. Given a set of input prompts, we compute the generated outputs and store the internal activations computed during the forward passes of the LM (e.g. the outputs of a specific Transformer [62] layer). The instances are then assigned a classification label meant to capture some property of the input and output. The probes are then trained to predict the label using only the internal activations as input. For example, Azaria and Mitchell [7] collect the activations of the middle layer of an LM when processing the output tokens, and manually label the generations as true or false. They then train a small neural network which can identify whether the model output is truthful given the internal activations of the generated tokens. Our work focuses on behaviors that can be trivially inferred from the output text alone (e.g. infer the answer from the output, infer if the output follows a specified format) and hence does not require manual labelling. Standard practice in probing uses the internal states of all tokens, even generated ones. We take an alternate approach, training probes on the internal states of the input tokens alone, allowing them to predict properties of the output tokens before they have been generated.
Conformal Prediction: Given a validation set, conformal prediction can be used to precisely calibrate the model’s confidence [17, 37]. During inference, the conformal system makes a prediction if and only if the probability of the prediction being accurate is above a user-specified probability [55].
In the classification setting, we are given a validation set $\{({\bf x_{1}},\ldots{\bf x_{n}}),(y_{1},\ldots y_{n})\},\mathbf{x}_{i}\in\mathbb{R}^{d},y_{i}\in\{1,2,\ldots,c\}$ (where $d$ is the dimension of the input vector and $c$ is the number of classes), a classifier which maps the input to scores for each class $f:\mathbb{R}^{d}\to\mathbb{R}^{c}$ and a user-defined confidence level $\alpha\in(0,1)$. To form a well-calibrated prediction set, we define a score function $S:\mathbb{R}^{d}\times\{1,2,\ldots c\}\to\mathbb{R}$ that measures how poor a prediction is e.g. $S(\mathbf{x},y)=1-(\frac{\exp{f(\mathbf{x})}}{\sum_{i}{\exp{f(\mathbf{x})}_{i}}})_{y}$. Using scores on the validation set $(s_{1},\ldots s_{n})$ we calibrate by calculating threshold $\hat{q}_{\alpha}$, the $1-\alpha$ quantile of the scores:
$\hat{q}_{\alpha}=\text{Quantile}(\{s_{1},s_{2},\ldots s_{n}\},\frac{\lceil(n+1)(1-\alpha)\rceil}{n})$ (1)
At inference time, the prediction set for a test instance $\mathbf{x}$ is simply $\{\hat{c}|S(\mathbf{x},\hat{c})\leq q_{\alpha}\}$.
For the score function provided above, no two classes can achieve a value greater than 0.5 simultaneously, implying that with $\alpha>0.5$ the prediction set at inference time is either empty (deferral/abstention) or consists of a single class label.
If the validation set is exchangeable (a slightly weaker assumption than the typical I.I.D assumption [8]) with the test set, then a prediction at inference time is provably guaranteed [55] to satisfy
$\mathbb{P}[y_{\text{test}}=\hat{c}_{\text{test}}]\geq 1-\alpha$ (2)
3 Can Internal States Predict Eventual Behavior?
Figure 2: Left: Estimation consistency (accuracy of predictions on the answer that CoT will eventually select. Right: Conformal estimation of CoT answer. Probes are often confident on sizeable portions of the test sets (37% on average) with low error even on datasets with poor total consistency.
In this section, we ask whether the internal states of Language Models contain information on the properties of the output tokens the model will eventually generate. We explain our methodology with the running example of a Chain-of-Thought prompted model answering a multiple choice question.
The CoT model is expected to reason about the input question before giving its final answer, a paradigm that has been shown to enable superior performance on a wide range of reasoning tasks [64, 36, 63]. Suppose we are given an input prompt of a question with two answer options:
   Question: Osteoclasts
   Option A: deposit bone and differentiate from monocytes.
   Option B: resorb bone and differentiate from monocytes.
   Explanation:
Llama3.1-8B [16] responds to this with
Osteoclasts are cells that break down bone tissue and are responsible for
bone reabsorption. They differentiate from monocytes and macrophages.
Answer: B
We can reasonably infer that the model will select option B by the time it generates the token ‘reabsorption’, but the tokens before that are generic enough that they do not eliminate either option. However, is it possible that the model contains signals that can identify the answer by the time it generates its initial output token ‘Osteoclasts’, and simply takes multiple tokens to express it in an interpretable way?
To test this, we collect 8 MCQ datasets—MMLU [22], CosmoQA [25], PiQA [9], ARC [13], MedMCQA [49], CommonsenseQA [59], OpenbookQA [44] and QASC [34] and use CoT to generate outputs with explanations before the answer (for more prompts see Appendix C).
We collect the output of the 18th (18/32) Transformer layer of Llama3.1-8B when processing the final token of the input (i.e. the final ‘:’), and train linear classifiers to predict what the eventual answer prediction will be. We ablate layers in Section 6 and reproduce results on other LMs in Appendix A.
The probes achieve (Figure 2) higher than random consistency (accuracy on prediction of model behavior) on all datasets, suggesting that input embeddings contain information on the eventual answer the LM will output. However, the performance varies significantly across datasets, suggesting that a naive attempt to exit early with these estimations could result in catastrophic failure.
3.1 Creating robust behavior estimators with conformal prediction
Unless one believes that the explanation generated by the LM has no causal influence on the eventual answer, there will always be instances where there is not enough information in the internal state of the input tokens to conclude what the future model behavior will be. The ideal behavior estimation system handles such cases, making consistent predictions when it is confident and deferring otherwise.
Hoping to impute such a capability to our probes, we use the probes we learned above in a conformal prediction framework. Specifically, we use a held-out validation set $\mathbb{D}_{\text{valid}}$ to calibrate the probe after training. We compute the probes’ prediction probabilities $\mathbf{\hat{y}}\in[0,1]^{|c|}$ for each class with true label $y\in\{1,\ldots c\}$ on the validation set, and find the lowest threshold quantile $q$ that satisfies:
$\frac{\sum_{\mathbf{y_{i}}\in\mathbb{D}_{\text{val}}}\mathbb{I}[(\max{(\mathbf{\hat{y_{i}}})\geq q)\land(\text{argmax}(\mathbf{\hat{y}_{i}})=\mathbf{y_{i}})}]}{\sum_{\mathbf{y_{i}}\in\mathbb{D}_{\text{valid}}}\mathbb{I}[\max{(\mathbf{\hat{y_{i}}})\geq q}]}\geq\alpha$ (3)
During inference, when the probe predicts a behavior with probability vector $\mathbf{\hat{y}}_{\text{test}}$, we return the prediction if and only if $\max{(\mathbf{\hat{y}}_{\text{test}})}\geq q$, otherwise we defer to the LM. In all experiments we set $\alpha=0.9$. If no satisfying $q$ can be found we defer on all test instances. We ablate $\alpha$ in Section 6.
Using the conformal probes (Figure 2) significantly increases the performance across all datasets. For datasets where the total test consistency was already high (ARC with $88\%$ total test consistency), the conformal probes predict on large portions of the test set with consistency close to 90%. On datasets where total test consistency was poor (MedMCQA, MMLU, PiQA), the probes predict on fewer instances. When they do predict, they do so with higher consistencies than on the full test set.
4 Accelerating Inference-Scaling with Conformal Probes
Figure 3: Tradeoff between accuracy loss (CoT Accuracy - Accelerated CoT Accuracy) and inference cost when using our method to accelerate CoT Prompting in IID (Left) and OOD (Right) settings. In IID setting, Accuracy loss is negative on average, and never more than 1.34%, while inference cost reductions are 65% on average. In the OOD setting, inference cost reductions are higher at 81% on average, with a small accuracy loss of 2.3% on average.
We extend the method described above to 18 other datasets spanning the tasks of Sentiment Analysis, Topic Analysis, Toxicity Detection and Fact Verification. Details on dataset setup are in Appendix C.
For all of these, we use Llama3.1-8B under CoT prompting (outputting an explanation before the final answer). We train a linear probe that uses the internal representation of the final input token at the 18th layer to predict the class that the CoT model will eventually output, and then perform conformal calibration. During inference, if the conformal probe is confident we use the probe estimation as the final answer. If not, we allow the model to continue its CoT generation and provide the final answer. We compare this to the vanilla CoT on two metrics—Accuracy Loss (Accuracy of CoT - Accuracy of Method) and Inference Cost Reduction (#CoT Forward Passes - #Method Forward Passes# CoT Forward $\frac{\#\text{CoT Forward Passes - \#Method Forward Passes}}{\text{\# CoT Forward Passes}}$).
The results (Figure 3) show that the method is highly effective at reducing the inference cost with minimal cost to accuracy. The minimum inference cost reduction is $4.7\%$, with an average reduction of $65\%$ across all datasets. Despite this significant speedup, the average accuracy loss is near zero ($\mathbf{-0.46\%}$), with the worst loss at $1.34\%$. Surprisingly, accuracy increases on several datasets.
Finally, we investigate the out-of-distribution generalization capabilities of the conformal probes. For each test dataset of the MCQ and Sentiment tasks, we train and calibrate using data from every other dataset. The results (Figure 3) show that the probes do exhibit OOD generalization, suggesting the method may be applicable even when there are slight shifts between training and test distributions.
5 Creating Early Warning Systems for Behaviors of Interest
(a) Unanswerable Question Abstention
(b) Bullet Point Format Following
(c) JSON Format Following
(d) Perplexity Confidence Estimation
(e) Verbalized Confidence Estimation
Figure 4: Probes can provide precise early warning signals for a range of important behaviours
The phenomenon of probes containing information on the eventual output sequence of the LM is not limited to predicting text classification answers but applies to a wide range of behaviors that can be represented with discrete labels. In this section, we showcase several other use cases of this finding. We describe the key experiment designs below, with exact prompts in the Appendix C.
LM Abstention: We collect two datasets (SelfAware [68], KnownUnknown [4]) which have both answerable and unanswerable questions. The LM is guided, via few-shot examples, to first reason out loud with respect to the question content and then either answer it or refuse to answer if the question seems to be unanswerable. Separately, we collect malicious and benign prompts from WildJailbreak [30], and prompt the LM to refuse to comply with a request if it is malicious. In both cases, the conformal probes try to estimate whether the LM will give an answer or refuse.
Results show that probes confidently estimate (Figure 4(a)) the abstention decision on over 15% of KnownUnknown, 44% of SelfAware and 83% of WildJailbreak, with over 90% across datasets.
Format Following Errors: We collect 3 QA datasets (NaturalQA [38], MSMarco [46], TriviaQA [32]) and prompt the LM to an answer in a specific format. One format requires the answer to be organized in exactly 3 bullet points, and the other requires a JSON output which contains pre-specified fields. The probes estimate whether the output will adhere to the format.
Predicting format following errors in a JSON format proves more challenging than the bullet point format( 4(b)). However, even on the JSON task, the probes maintain a minimum consistency of 84% and have confidence on over 8% of instances on 2/3 datasets while deferring completely on the other.
Low Confidence Output: On the same 3 QA datasets, we try to identify when the LM is less confident in its response. Following recent work [33, 31], we consider one setting where we explicitly prompt the model to verbalize its confidence, and another where we use the per-token perplexity of the output as a proxy for the confidence. In the case of the perplexity measure, we consider the bottom $25\%$ of scores to be ‘high confidence’, and the top $25\%$ to be ‘low confidence’ (discarding the rest). In both cases, the probes attempt to identify whether the model will be confident in its output.
This is the most challenging task, probes typically make predictions on around 4% of test instances. However, consistency remains high, making the early warning system feasible in practice.
The conformal probe-based early warning systems vary in their confidence across tasks and datasets. However, on a wide range of LM behaviors, they predict on sizable portions of the test set with consistently high consistency.
6 Analysis and Ablations
Figure 5: Left: Density plot of the correlation coefficient between the CoT model accuracy and the consistency of the probe estimate. The coefficient is often positive, indicating that instances where the probe inaccurately estimates CoT model are also instances where the model is more likely to be incorrect. Right: On CosmoQA, probes have near-random consistency with initial tokens. There is a sudden performance boost towards the final token. Confidence remains steady, suggesting poor calibration in early token probes.
Why does accuracy improve when using probes? Seeking to explain the surprising increase in accuracy when accelerating CoT models (Figure 3), we plot (Figure 5) the correlation between the CoT models accuracy and the probes estimation consistency. It is generally a positive correlation, suggesting that on the instances where the probe estimates incorrectly, the CoT model is also more likely to output an incorrect answer. This minimizes the harm of an inconsistent estimation as sometimes making an inaccurate estimate is actively beneficial for overall accuracy.
How early in the computation does the model show signs of the final behavior? We have shown that the final input token often contains sufficient information to predict output behavior, but at which token does this information start becoming clear and accessible? To investigate this, we collect the activations of the mid-layer from every input token (after the FewShot example tokens) and train a linear classifier to map any of these internal activations to the desired output behavior.
Probing (Figure 5) using the embeddings of tokens in the first quarter of the question tokens leads to near-random performance. While consistency increases steadily as we use later tokens, a considerable jump is noticed when shifting to the final tokens. Surprisingly, confidence declines marginally as we use later tokens for probing. This suggests that using the earliest tokens for probing could result in inconsistent and poorly calibrated models, making it a poor design decision for this method.
How does output length affect estimation accuracy? We measure the correlation between the estimation consistency of the probe and the token count of both the input (-0.41%) and output (-6.1%). While input length seems to have little effect, a higher output length is correlated with worse accuracy. This suggests probes struggle more with inputs that evoke longer outputs.
Figure 6: Probes achieve high accuracy with as few as 1000 datapoints in most cases
How much data does a probe need? We ablate the amount of data used to train the probe. We find (Figure 6) that often, fewer than 1000 datapoints is sufficient for the probe to achieve high estimation accuracy, suggesting that the method is particularly data efficient.
Figure 7: Layer ablations show that early layers perform poorly, while the decision between mid and late layers is dataset specific
Which layer should we use to probe? We explore (Figure 7) the specific layer being probed affected the performance of the probe. The most general finding is that the early layers offer poor estimation consistency. However, whether mid or late layers perform better is task and dataset-specific, and cannot be described generally. For more details see Appendix B.
Figure 8: Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero
How does varying $\alpha$ change performance? Increasing the confidence required from the probes has predictable effects (Figure 8). Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
Figure 9: Probe accuracy increases with model size, suggesting the method may scale favorably
How does model scale affect performance? We varied the size of the LM used (Llama3 3B, 8B and 70B) and measure probe consistency across several layers. Encouragingly (Figure 9), the performance of the larger models are consistently better, suggesting that the information encoded in the internal activations is easier for the probe to ‘read’ when the model is more powerful. This bodes well for the methods’ ability to scale with models of increasing size and capability.
What are the limits of the probes? Using the MCQ task, we tried to estimate whether the LM would output an incorrect answer. The probe accuracies were consistently poor and near random. This shows that, intuitively, the probes have limitations in the behaviors they can estimate. We hypothesize that since probes are simple linear classifiers, they can only detect patterns and use ‘knowledge’ that is well encoded in the internal activations being probed. This suggests that probes struggle with output properties that cannot be identified by the LM itself (without external knowledge).
7 Related Works
This work takes inspiration from recent methods that probe the hidden states of LMs to observe interpretable patterns [3, 35, 50, 23, 2] identify false statements [7, 39, 40, 69] and hallucinations [12, 57, 29]. Our work also has connections to literature on early exiting during the forward pass of NN models, with works often using signals from the hidden states to prematurely exit with a prediction on the next token [65, 70, 66, 54, 27]. Notably,  Schuster et al. [54] also uses a conformal prediction framework to give a provable error bound on the next token approximation.
We are the first to show that the internal states can predict a range of behaviors, uninferable from the next token alone, before any of the output tokens are generated. Additionally, we are the first to show that conformal prediction can be used to create early warning systems for a wide range of behaviors like question abstention to format following errors. Our work advances research on understanding the nature of the information contained in the hidden states of LMs [50, 5, 47, 41, 61]. Specifically, we show that the information contained in the hidden states is relevant not just to the next token, but to behaviors that manifest several tokens later during the LMs generation.
8 Conclusion
We show that a Language Model’s internal representation of input tokens alone contains vital information on the behavior of the LM over the entire output sequence. These signals can be used to create probes that serve as precise early warning or exit systems for LMs on a wide range of behaviors. On 27 text classification datasets across 5 different tasks, the method can accelerate Chain-of-Thought prompting by 65% with little accuracy loss. The method can be used to create high-precision early warning systems that identify if an LM will refuse to answer a question, fall victim to jailbreaking, fail to follow output format specifications or give low-confidence responses. We show that the probes generalize to out-of-distribution test sets, and scale favorably to larger LMs. Finally, we explore the limitations of the method, showing that the behavior of longer output sequences is harder to estimate and that tasks that require knowledge external to the model are particularly challenging. With the rising popularity of inference-time scaling methods, we hope our work can help ameliorate the growing computational cost of running LMs and provide more insight into the nature of the information contained in the hidden states of LMs.
Appendix A Confirming Robustness of Results
A.1 Confirming Robustness to Choice of Model
We conduct experiments on two other Language Models, to ensure that our results hold on Language Model families outside of the Llama3 series. We use Mistral-7B-Instruct-v0.3 from MistralAI [28] and DeepSeek-R1-Distill-Qwen-14B from DeepSeek [20]. The results show that the phenomenon shown in the main text can be detected robustly across various model families.
Figure 10: Deepseek
Figure 11: Mistral
A.2 Data Leakage
Several datasets used in this study have been released before the creation of Llama3, and as such might have been pre-exposed to the model during training. We detail our reasons as to why data leakage is not a concern for the conclusions that we draw from this study:
  • There are datasets in our study that have not been leaked. For example, the MMLU test set is used as an official test benchmark by LLama3[16] and WildJailbreak was released to the public only after the initial release of Llama3 models [30]. The results are strong on such datasets as well, showing that data leakage is not behind the performance of the probes.
  • Several tasks require the probes to identify behavior that fundamentally does not exist in text form on the internet. The bullets, JSON and Confidence tasks require the probes to determine how the model will behave, regardless of whether the model has seen the input text or not. For example, whether or not the model was pretrained on MSMarco questions is immaterial for predicting whether or not it will output answers in a specific JSON format.
Appendix B Extended Analysis
B.1 Layer Ablation
(a) Immediate saturation
(b) Only confidence rises and then plateaus
(c) Confidence rises and dips, accuracy rises past middle
(d) Confidence rises and dips, accuracy dips after middle
(e) Accuracy stays roughly constant, confidence rises slightly then falls back to starting
(f) Accuracy constant, significant variance in confidence with peak in middle
Figure 12: Layer ablations show that the decision on which layer to use is highly dependant on the specific task and dataset.
A deeper analysis of the interaction between the performance and the layer used for probing shows that no consistent patterns can be identified. A subset of the common patterns (Figure 12) shows that sometimes that apart from early layers being poor, it is unclear whether the middle or final layers are consistently better for accuracy or confidence.
Appendix C Experiment Details
C.1 Datasets:
We used the following datasets in our experiments, all usage is in accordance with their respective licenses. For each dataset, we select a maximum of 50,000 training instances to train (and validate) our probes, using the full test set to measure all metrics.
C.1.1 Multiple Choice Question Answering:
To test this, we collect 8 MCQ datasets—MMLU [22], CosmoQA [25], PiQA [9], ARC [13], MedMCQA [49], CommonsenseQA [59], OpenbookQA [44] and QASC [34] and use CoT to generate outputs with explanations before the answer (for more prompts see Appendix C).
ARC: The AI2 Reasoning Challenge (ARC) [13] is a knowledge and reasoning challenge that contains 7,787 natural, grade-school science questions (authored for human tests).
CommonsenseQA: The CommonsenseQA [59] dataset contains 12,247 questions with complex semantics that often require prior knowledge.
MedMCQA: A large-scale, Multiple-Choice Question Answering (MCQA) dataset designed to address realworld medical entrance exam questions, MedMCQA [49] has 194k AIIMS and NEET PG entrance exam MCQs covering 2.4k healthcare topics and 21 medical subjects.
MMLU: The Massive Multitask Language Understanding Benchmark [22] is a test to measure a text model’s multitask accuracy. The test covers 57 tasks including elementary mathematics, US history, computer science, law, and more.
OpenBookQA: Modeled after an open book examination, OpenBOOkQA [44], OpenBookQA consists of 6000 questions that probe an understanding of elementary level science facts and their application to novel situations.
QASC: Question Answering via Sentence Composition [34] tests a models ability to compose knowledge from multiple pieces of texts. The dataset consists of 9,980 multiple-choice questions from elementary and middle school level science, with a focus on fact composition;
PiQA: Physical Interaction: Question Answering [9] is a dataset that tests whether AI systems can learn to reliably answer physical common-sense questions without experiencing the physical world. The dataset consists of 16,000 training QA pairs with 2,000 and 3,000 examples held out for validation and training.
CosmoQA: Commonsense Machine Comprehension Question Answering [25] is a dataset of 35,600 problems that require commonsense-based reading comprehension, formulated as multiple-choice questions. The dataset focuses on reading between the lines with questions that require reasoning beyond the exact text spans in the context.
C.1.2 Sentiment Analysis:
MTEB: The following datasets were taken from the Massive Text Embedding Benchmark [45]. AmazonReviews: a dataset with 1.7 million Amazon product reviews and a sentiment score ranging from 0-5, TwitterSentiment: a dataset with 30,000 tweets and a sentiment class from positive, neutral or negative
Yelp Polarity: The Yelp review rating challenge [6] consists of nearly 600,000 yelp reviews with sentiment classes from positive or negative.
TwitterFinance: The Twitterfinance dataset [1] is an annotated corpus of 11,000 finance-related tweets. This dataset is used to classify whether the tweet are bullish, bearish, or neutral.
NewsMTC: A dataset for sentiment analysis (TSC) on news articles reporting on policy issues, NewsMTC [21] consists of more than 11,000 labeled sentences.
IMDB Reviews: A classic sentiment analysis dataset, IMDB Reviews [42] consists of 50,000 highly polar movie reviews with binary labels.
Financial Phrasebank: A dataset that measures the polar sentiment  [43] of sentences from financial news. The dataset consists of 4840 sentences from English language financial news categorised by sentiment. The dataset is divided by agreement rate of 5-8 annotators.
AuditorSentiment: Based on Financial Phrasebank, this dataset [26] is additionally annotated by auditors to reflect bearish, bullish and neutral labels for accounting related sentences.
Emotion: The DAIR-AI Emotion dataset [52] is a dataset of 20,000 Twitter messages with six basic emotions: anger, fear, joy, love, sadness, and surprise.
SST5: A standard sentiment analysis dataset, SST5 [56] consists of fine grained sentiment labels for 215,154 phrases in the parse trees of 11,855 sentences.
C.1.3 Fact verification:
ClimateFEVER: A dataset that consists of 1,535 real-world claims regarding climate-change collected on the internet. Each claim In ClimateFEVER [15]is accompanied by five manually annotated evidence sentences retrieved from the English Wikipedia that support, refute or do not give enough information to validate the claim totalling in 7,675 claim-evidence pairs.
HealthVER: A dataset for evidence-based fact-checking of health-related claims, HealthVER [53] consists of 14,330 evidence-claim pairs.
FEVER: Fact Extraction and VERification [60] consists of 185,445 claims generated by altering sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence they were derived from.
C.1.4 Topic Identification:
AGNews: A collection of more than 1 million news articles. AGNews articles have been gathered from more than 2000 news sources and annotated for Topic [19].
BBCNews: The BBCNews dataset [18] is a dataset consisting of 2,225 articles published on the BBC News website corresponding during 2004-2005. Each article is labeled under one of 5 categories: business, entertainment, politics, sport or tech.
NYTimes: The New York Times Annotated Corpus [51] contains over 1.8 million articles written and published by the New York Times between January 1, 1987 and June 19, 2007 with article metadata provided by the New York Times Newsroom, the New York Times Indexing Service and the online production staff at nytimes.com.
C.1.5 Toxicity Detection:
We use two datasets provided by the same organization, JigsawToxicity  and a JigsawUnintendedBiasToxicity . Both datasets have scores for toxicity from a chatroom setting.
C.1.6 Others:
Unanswerable Questions: We use two datasets which contain unanswerable questions, SelfAware [68] and KnownUnkown [4]. Selfaware is a dataset consisting of unanswerable questions from five diverse categories and their answerable counterparts. KnownUnknown is a dataset with questions on known quantities and unknown quantities. We collect jailbreaking prompts (and benign prompts) from WildJailbreak [30], a dataset with 262,000 vanilla (direct harmful requests) and adversarial (complex adversarial jailbreaks) prompt-response pairs. The dataset also has benign prompts that should be complied with.
Format Following, Confidence Elicitation: Both of these tasks use the same set of datasets - NaturalQA [38], TriviaQA [32] and MSMarco [46]. These datasets were selected as they are large question-answering datasets which often require generative output that can vary in length.
C.2 Task Specific Set Up:
C.2.1 Text Classification Tasks:
In all text-classification tasks, the CoT model is first asked to output a reasoning chain and then finally provide a classification answer. The probes are always trained to preemptively identify what the prediction will be using only the input token embeddings.
MCQ: For the MCQ tasks, we select two options by randomly sampling an incorrect answer from the provided options along with the correct option. This is because the MCQ datasets vary in terms of number of MCQ options, and for the out-of-distribution experiment we require all the datasets to have the same number of answer classes. The CoT model is asked to identify the correct answer option.
Sentiment: For the MCQ task, many datasets are in a binarized format, while others include continuous scores, or ternary labels (including neutral sentiment. We make all the datasets have a similar label structure, by keeping only positive and negative labels. Concretely, the label is 1 if and only if the sentiment is positive (or bullish for finance datasets), and 0 otherwise. The CoT model is asked to identify whether or not the text has a positive sentiment.
Topic: We keep only a subset of the topics to ensure every topic is fairly represented in the data. After dropping topics, AGNews has World, Sports, Businsess or Science, BBCNews has Business, Tech, Sports or Politics, while NYTimes has Health, Fashion, Real Estate or Television. The CoT model is asked to identify the topic of the article.
Fact Verification: We map all claims to with supported or not supported (includes refuted and neutral). For ClimateFEVER and HealthVer, we provide evidence along with the claim, while for FEVER we provide only the claim. The CoT model is asked to identify whether or not the claim is supported.
Toxicity Detection: We randomly sample the datasets to ensure a balance of toxic and not toxic comments. The CoT model is asked to detect whether or not the text is toxic
C.2.2 Other Tasks:
In these tasks, the probes are trained to predict a variety of other behaviors using the input token embeddings.
LM Abstention: For the unanswerable question datasets of SelfAware and KnownUnknown, the model is first instructed to give a reasoning chain about the question, and answer only if the question is answerable (abstaining otherwise). With WildJailbreak, the input need not be a question, but any request. Hence, we instruct the model to comply with the request if it is not malicious, and abstain otherwise. In all cases, the probe is trained to detect whether or not the LM will abstain.
Confidence Estimation: There are two settings for this task:
  • Internal Confidence: The LM is prompted to output an answer to the input question, and we record the token-normalized perplexity as a proxy of confidence. We keep only the bottom and top 25% of instances as per normalized perplexity and train the probes to differentiate between the two.
  • Verbal Confidence: The LM is prompted to output an answer, along with a confidence score (either confident or unsure). The probe is trained to identify what the confidence score will be.
Format Following: Using an input prompt, we specify two different output formats a LM must obey. In both cases, the probe learns whether the LM will fail to follow format specifications or comply with them.
  • Bullets: The output should be presented in 3 numbered bullet points, no fewer and no more
  • JSON: The output should be presented in a JSON string with the following structure: {[‘short_answer’]: <str>, ‘entities’: List<str>, ‘references’: List<str>}
C.3 Prompts:
We have written separate prompts for each dataset to ensure the LM follows the instructions and outputs the text in a way that guarantees we can parse it and infer the behavior we seek to pre-emptively identify with the probes. For example, a prompt for the MMLU dataset for MCQA is :
 Question: What is true for a type-Ia supernova?
 Option A: This type occurs in young galaxies
 Option B: This type occurs in binary systems
 Give an explanation and then the answer:
 Explanation: Type Ia supernova is a type of supernova that occurs when
 two stars orbit one another in which one of the stars is a white dwarf
 Answer: B
 Question: <NEW QUESTION>
 Give an explanation and then the answer:
 Explanation:
For the JSON task, one such prompt is:
Answer the following questions by giving a short_answer, entities list
 and references list. Give the output in JSON format
 Question: What is the capital of France?
 Answer: { "short_answer": "Paris",
 "entities": ["France"], "references": ["https://en.wikipedia.org/wiki/Paris"]}
 Question: <NEW QUESTION>
 Answer:
Each task and dataset has a different set of prompts, we have provided a link to the code https://github.com/DhananjayAshok/LMBehaviorEstimation. All prompts can be seen in the file data.py
C.4 Hardware:
All of our experiments were run on a compute cluster with 8 NVIDIA A40 GPUs (approx 46068 MiB of memory) on CUDA version 12.6. The CPU on the cluster is an AMD EPYC 7502 32-Core Processor. Most experiments could be conducted with less than 16GB of RAM.
C.5 Replicability:
To ensure that our code is easy to replicate and our method is easy to extend, we have provided open access to our code at https://github.com/DhananjayAshok/LMBehaviorEstimation

### v2

@@
 Abstract
-Autoregressive Language Models output text by sequentially predicting the next token to generate, with modern methods like Chain-of-Thought (CoT) prompting achieving state-of-the-art reasoning capabilities by scaling the number of generated tokens.
-However, are there times when we can infer how the model will behave (e.g.
-abstain from answering a question) early in the computation, making generation unnecessary?
-We show that internal representation of input tokens alone can often precisely predict, not just the next token, but eventual behavior over the entire output sequence.
-We leverage this capacity and learn probes on internal states to create early warning (and exit) systems.
-Specifically, if the probes can confidently estimate the way the LM is going to behave, then the system will avoid generating tokens altogether and return the estimated behavior instead.
-On 27 text classification datasets spanning five different tasks, we apply this method to estimate the eventual answer of an LM under CoT prompting, reducing inference costs by 65% (average) while suffering an accuracy loss of no more than 1.4% (worst case).
-We demonstrate the potential of this method to pre-emptively identify when a model will abstain from answering a question, fail to follow output format specifications, or give a low-confidence response.
-We explore the limits of this capability, showing that probes generalize to unseen datasets, but perform worse when LM outputs are longer and struggle to predict properties that require access to knowledge that the models themselves lack.
-Encouragingly, performance scales with model size, suggesting applicability to the largest of models.
-[^1] https://github.com/DhananjayAshok/LMBehaviorEstimation
+The text produced by language models (LMs) can exhibit specific ‘behaviors,’ such as a failure to follow alignment training, that we hope to detect and react to during deployment.
+Identifying these behaviors can often only be done post facto, i.e., after the entire text of the output has been generated.
+We provide evidence that there are times when we can predict how an LM will behave early in computation, before even a single token is generated.
+We show that probes trained on the internal representation of input tokens alone can predict a wide range of eventual behaviors over the entire output sequence.
+Using methods from conformal prediction, we provide provable bounds on the estimation error of our probes, creating precise early warning systems for these behaviors.
+The conformal probes can identify instances that will trigger alignment failures (jailbreaking) and instruction-following failures, without requiring a single token to be generated.
+An early warning system built on the probes reduces jailbreaking by 91%.
+Our probes also show promise in pre-emptively estimating how confident the model will be in its response, a behavior that cannot be detected using the output text alone.
+Conformal probes can preemptively estimate the final prediction of an LM that uses Chain-of-Thought (CoT) prompting, hence accelerating inference.
+When applied to an LM that uses CoT to perform text classification, the probes drastically reduce inference costs (65% on average across 27 datasets), with negligible accuracy loss.
+Encouragingly, probes generalize to unseen datasets and perform better on larger models, suggesting applicability to the largest of models in real-world settings.
+[^1] Our code is accessible at: https://github.com/DhananjayAshok/LMBehaviorEstimation
 1 Introduction
-Language Models trained to predict the next token of a sentence on large-scale data [58, 14] are unrivalled in general Natural Language Understanding (NLU).
-During inference, these models generate remarkably high-quality outputs by sequentially predicting the most appropriate output tokens to generate.
-Frontier LM systems of today build on this by scaling inference i.e.
-increasing the number of tokens or sequences generated [64, 67, 48, 20], to provide even better responses.
-In this paper we ask whether LMs have learned representations that contain information, not just about the next token the model will predict, but the entire sequence as a whole.
-We conclude with an affirmative answer, showing that the internal states of the LM can often be used to preemptively identify how an LM will behave.
-This information becomes accessible before the LM has generated a single token, potentially obviating the need for further inference on those instances.
-Concretely (Figure 1), we train linear classifiers (probes) [3] that use the internal representation of the input tokens to predict the eventual behavior of the LM (e.g.
-whether an LM will comply with a request or abstain).
-We calibrate the probes using methods from conformal prediction [55], allowing them to estimate behavior only when confident.
-During inference, probes predict only when there is a provable guarantee on the estimation error, ensuring precise early warning signals for various model behaviors.
-We apply this method to an LM that tackles text classification by generating several explanation tokens before a class prediction (CoT).
-The probes estimate the final CoT prediction and exit early if confident in their estimate.
-On 27 datasets, spanning the tasks of MCQA, Sentiment Analysis, Topic Classification, Toxicity Detection and Fact Verification, our method reduces inference costs by 65% on average, while suffering accuracy losses of no more than 1.4%.
-Furthermore, on 14/27 of the datasets, our method reduces inference costs by 63% on average with no cost to accuracy.
+Language models (LMs) have emerged as the dominant approach to general language tasks [88, 22], seeing widespread adoption in chatbots [3, 68], code generation [18] and reasoning systems [67].
+However, they have been known to ‘misbehave,’ hallucinating false information [98], failing to adhere to output specifications [59], or, most concerningly, producing content that is misaligned with human values [75], sometimes dangerously so [35].
+In practice, such misbehavior is mitigated by deploying post-hoc guardrails on the output of the LM [92, 24, 16, 101, 8, 53].
+For example, a chatbot LM may exhibit harmful behavior by complying with a malicious request [41, 90]; developers will typically attempt to detect this behavior in the output [101] and return an abstention message instead of the original LM content.
+Such a framework suffers the economic [100] and environmental [83, 76] expense of generating every output token, a cost that threatens to grow exponentially as frontier labs continue to scale both model sizes [37, 49] and the number of tokens generated during inference [91, 95, 67, 32].
+In this paper, we show that the internal representations of LMs can be used to preemptively predict behaviors that will emerge over the entire output sequence.
+This information becomes accessible before the LM has generated a single token, enabling the system to take appropriate actions without suffering inference costs on those instances.
+Concretely (Figure 1), we train linear classifiers (probes) [4] that use an LM’s internal representation of input tokens to predict the eventual behavior of its output.
+We then calibrate the probes using methods from conformal prediction [81].
+During inference, we look to probes to make a prediction only when there is a provable guarantee on the estimation error, ensuring precise early warning signals for various model behaviors.
+We show that these probes can preemptively identify degenerate behavior, such as instruction-following failures (Section 3).
+Across multiple datasets and output formats, our probes give precise early warning signals for cases where the LM will fail to follow specified output format instructions.
+Conformal probes can also detect deeper, ‘self-reflective’ properties, like whether an LM’s behavior conflicts with its safety alignment.
+We demonstrate this (Section 4.1) by applying our method to enhance the safety of an LM that is instructed to abstain from complying with malicious requests.
+Probes are trained to identify cases where this will not happen—the request will be malicious and the model will not abstain.
+Once again, our probes are able to precisely detect (with over 92% accuracy) cases where the LM complies with malicious requests.
+An early warning system built on these conformal probes reduces the jailbreak success rate from 30% to 2.7%, a 91% reduction.
+We further show that conformal probes can identify behaviors that cannot be measured with output text.
+We show (Section  4.2) that our probes can a priori estimate an LM’s confidence as measured by the per-token perplexity of the output; as this model-specific property cannot be inferred from just the text sequence alone, we show the wide scope of applicability of our probing approach.
+Finally, we apply our method to an LM that tackles text classification by generating several explanation tokens before a class prediction, i.e.
+using Chain-of-Thought (CoT) reasoning [91] (Section 5).
+The probes estimate the final class prediction that will appear after the reasoning chain and exit early if confident in their estimate.
+On 27 datasets, spanning the tasks of Multiple Choice QA, Sentiment Analysis, Topic Classification, Toxicity Detection and Fact Verification, our method reduces inference costs by 65% on average, while suffering accuracy losses of no more than 1.4% (worst case).
 The probes generalize to unseen datasets, reducing the inference cost of CoT on OpenBookQA [63] by 68% with minimal loss to accuracy (0.4%), despite only training on other MCQA datasets.
-Beyond accelerating text classification, probes can pre-emptive identify several behaviors of interest.
-We apply our method to a QA system that verbalizes some reasoning about the input question and then either answers or abstains if the question is deemed to be unanswerable.
-The probes, which are trained to estimate whether the LM will abstain, make confident predictions on over 15% of the test instances and achieve an estimation consistency of over 90% when doing so.
-Probes can provide precise early warning signals for degenerate behavior (whether the model will fail to follow a specified output format) and even ‘self-reflective" properties which cannot be inferred from the text output alone (whether the model will output a high perplexity answer to a question).
-While the number of instances where the probes give confident predictions varies significantly across datasets and tasks, the probe maintains high consistency on confident predictions.
-We show that often fewer than a thousand training instances are sufficient for the probes to attain high estimation consistency, while ablations on the probing layer present a more complicated, task-specific story.
-Encouragingly, increasing the scale of the LM improves the performance of the probe, suggesting the method may scale favorably with ever-increasing model sizes [10, 24, 11, 16].
-The probes are limited in the behaviors they can preemptively detect.
-While probes can identify various model behaviors, they fail on those that cannot be identified without knowledge external to the model (e.g.
-identifying whether the model is likely to output an incorrect answer to an MCQ instance).
-Additionally, probes are less confident on instances which trigger longer outputs.
+We demonstrate that often fewer than 500 training instances are sufficient for the probes to attain high estimation consistency, while ablations on the probing layer present a more complicated, task-specific story.
+Encouragingly, increasing the scale of the LM improves the performance of our technique, suggesting it may scale favorably with ever-increasing model sizes [15, 37, 19, 26].
 We hope our findings enable practitioners to build efficient early warning systems on language models and enable further research into the information encoded in their hidden states [71, 10, 66].
-Figure 1: Overview of our method: During training we collect the hidden states of the input token and eventual output behavior on a variety of prompts, using them to train a probe on the hidden states.
+Figure 1: Overview of our method as applied in preemptive safety alignment failure detection.
+During training, we use many input prompts to collect a dataset consisting of the hidden states of the input tokens and eventual output behavior (in this example, the behavior is whether the LM fails to abstain on a malicious prompt).
+The dataset is used to train a probe on the hidden states.
 With a validation set, we calibrate the probes for conformal prediction by identifying a threshold for confident predictions.
-During inference, if the probes are confident in their estimate of the eventual LM behavior, we stop all computation and return the prediction.
+During inference, if the probe is confident in its estimate of the eventual LM behavior, we stop computation early (in this example, overriding the LM and abstaining).
 2 Background
 Hidden State Probing: Lightweight probes have long been used to interpret the internal activations of neural networks [4] and language models [71, 10].
-Given a set of input prompts, we compute the generated outputs and store the internal activations computed during the forward passes of the LM (e.g.
+Given a set of input prompts, we compute the generated outputs and store the internal activations computed during the forward passes of an LM (e.g.
 the outputs of a specific Transformer [88] layer).
@@
 For example, Azaria and Mitchell [10] collect the activations of the middle layer of an LM when processing the output tokens, and manually label the generations as true or false.
-They then train a small neural network which can identify whether the model output is truthful given the internal activations of the generated tokens.
-Our work focuses on behaviors that can be trivially inferred from the output text alone (e.g.
-infer the answer from the output, infer if the output follows a specified format) and hence does not require manual labelling.
-Standard practice in probing uses the internal states of all tokens, even generated ones.
-We take an alternate approach, training probes on the internal states of the input tokens alone, allowing them to predict properties of the output tokens before they have been generated.
+They then train a small neural network which can identify whether the model output is truthful, given the internal activations of the generated tokens.
+Prior work in probing uses the internal states of all tokens, even generated ones, which incurs the cost of inference.
+In this work, we explore an alternate approach, training probes on the internal states of the input tokens alone, allowing us to predict properties of the output tokens before they have been generated.
 Conformal Prediction: Given a validation set, conformal prediction can be used to precisely calibrate the model’s confidence [29, 52].
@@
 $\hat{q}_{\alpha}=\text{Quantile}(\{s_{1},s_{2},\ldots s_{n}\},\frac{\lceil(n+1)(1-\alpha)\rceil}{n})$ (1)
-At inference time, the prediction set for a test instance $\mathbf{x}$ is simply $\{\hat{c}|S(\mathbf{x},\hat{c})\leq q_{\alpha}\}$.
-For the score function provided above, no two classes can achieve a value greater than 0.5 simultaneously, implying that with $\alpha>0.5$ the prediction set at inference time is either empty (deferral/abstention) or consists of a single class label.
+At inference time, the prediction set for a test instance $\mathbf{x}$ is $\{\hat{c}|S(\mathbf{x},\hat{c})\leq q_{\alpha}\}$.
+For the above score function, no two classes can score greater than 0.5 simultaneously, implying that with $\alpha>0.5$ the prediction set at inference time is either empty (deferral/abstention) or consists of a single class label.
 If the validation set is exchangeable (a slightly weaker assumption than the typical I.I.D assumption [12]) with the test set, then a prediction at inference time is provably guaranteed [81] to satisfy
@@
 3 Can Internal States Predict Eventual Behavior?
-Figure 2: Left: Estimation consistency (accuracy of predictions on the answer that CoT will eventually select.
-Right: Conformal estimation of CoT answer.
-Probes are often confident on sizeable portions of the test sets (37% on average) with low error even on datasets with poor total consistency.
-In this section, we ask whether the internal states of Language Models contain information on the properties of the output tokens the model will eventually generate.
-We explain our methodology with the running example of a Chain-of-Thought prompted model answering a multiple choice question.
-The CoT model is expected to reason about the input question before giving its final answer, a paradigm that has been shown to enable superior performance on a wide range of reasoning tasks [64, 36, 63].
-Suppose we are given an input prompt of a question with two answer options:
-   Question: Osteoclasts
-   Option A: deposit bone and differentiate from monocytes.
-   Option B: resorb bone and differentiate from monocytes.
-   Explanation:
-Llama3.1-8B [16] responds to this with
-Osteoclasts are cells that break down bone tissue and are responsible for
-bone reabsorption. They differentiate from monocytes and macrophages.
-Answer: B
-We can reasonably infer that the model will select option B by the time it generates the token ‘reabsorption’, but the tokens before that are generic enough that they do not eliminate either option.
-However, is it possible that the model contains signals that can identify the answer by the time it generates its initial output token ‘Osteoclasts’, and simply takes multiple tokens to express it in an interpretable way?
-To test this, we collect 8 MCQ datasets—MMLU [22], CosmoQA [25], PiQA [9], ARC [13], MedMCQA [49], CommonsenseQA [59], OpenbookQA [44] and QASC [34] and use CoT to generate outputs with explanations before the answer (for more prompts see Appendix C).
-We collect the output of the 18th (18/32) Transformer layer of Llama3.1-8B when processing the final token of the input (i.e.
-the final ‘:’), and train linear classifiers to predict what the eventual answer prediction will be.
-We ablate layers in Section 6 and reproduce results on other LMs in Appendix A.
-The probes achieve (Figure 2) higher than random consistency (accuracy on prediction of model behavior) on all datasets, suggesting that input embeddings contain information on the eventual answer the LM will output.
-However, the performance varies significantly across datasets, suggesting that a naive attempt to exit early with these estimations could result in catastrophic failure.
+In this section, we ask whether the internal representations of the LMs input can be used to predict eventual behavior, even before any output token has been generated.
+We explain our methodology with the running example of an LM instructed to answer a question in bullet point format.
+The LM is expected to provide an answer in two bullet points, with any fewer or greater bullets in the answer being a failure to follow instructions. Suppose we are given an input prompt:
+   Question: What do Osteoclasts do?
+A LM may respond to this with:
+1. Decompose bone 2. Differentiate from monocyte 3. Create space for tissue
+We can identify that the LM has behaved contrary to its instructions when it generates the ‘3’.
+The previous tokens give no indication that a failure was imminent; however, is it possible that the model contains signals that can identify this eventual failure by the time it generates the first token, i.e.
+‘1’?
+[…]
+To test this, we collect three QA datasets—NaturalQA [54], MSMarco [65] and TriviaQA [47] and prompt the LM to answer questions in a specific format.
+One format requires the answer to be organized in exactly three bullet points, while the other requires a JSON output with pre-specified fields.
+We evaluate whether the output format has been followed, and then sample this data to obtain training and testing splits that have an equal number of failures and successes.
+We collect the output of the middle Transformer layer of Llama3.1-8B [26] when processing the final token of the input (in the example above, the ‘?’), then train linear classifiers to predict whether the output will eventually fail to follow instructions.
+We ablate layers in Section 6, reproduce results on other LMs in Appendix A, and provide the prompts that specify the format in Appendix C.
+We also fine-tune a bert-large-uncased [22] model for text classification, which involves tuning 340M parameters, $4\times 10^{4}$ times the number of trainable parameters of the linear model.
+The BERT model is trained to predict whether the output will fail to follow formatting instructions given the text of the input prompt alone, and is a measure of how a significantly stronger model performs when attempting to establish a pattern between the inputs to the model and the output behavior.
+Methods are evaluated on the metric of estimation consistency, i.e, how accurate they are at preemptively predicting whether or not the formatting instructions will be followed.
+The probes (Figure 2, navy blue bar) outperform the random baseline in all settings, showing that an LM’s representation of the input tokens contains information on behavior that will emerge over the entire output sequence.
+Moreover, the probes also outperform fine-tuning on BERT in all but one of the settings, despite using around only 0.0025% as many parameters.
+However, despite outperforming the baselines, the average estimation consistency of the probes is less than 75%. A naive attempt to act on the probes estimate would result in catastrophic failure.
 3.1 Creating robust behavior estimators with conformal prediction
-Unless one believes that the explanation generated by the LM has no causal influence on the eventual answer, there will always be instances where there is not enough information in the internal state of the input tokens to conclude what the future model behavior will be.
-The ideal behavior estimation system handles such cases, making consistent predictions when it is confident and deferring otherwise.
-Hoping to impute such a capability to our probes, we use the probes we learned above in a conformal prediction framework.
+Internal states may occasionally be insufficient to estimate eventual behavior.
+The ideal system handles such cases, making consistent predictions when confident and deferring otherwise.
+Hoping to impart such a capability to our probes, we use the probes learned above in a conformal prediction framework.
 Specifically, we use a held-out validation set $\mathbb{D}_{\text{valid}}$ to calibrate the probe after training.
@@
 $\frac{\sum_{\mathbf{y_{i}}\in\mathbb{D}_{\text{val}}}\mathbb{I}[(\max{(\mathbf{\hat{y_{i}}})\geq q)\land(\text{argmax}(\mathbf{\hat{y}_{i}})=\mathbf{y_{i}})}]}{\sum_{\mathbf{y_{i}}\in\mathbb{D}_{\text{valid}}}\mathbb{I}[\max{(\mathbf{\hat{y_{i}}})\geq q}]}\geq\alpha$ (3)
-During inference, when the probe predicts a behavior with probability vector $\mathbf{\hat{y}}_{\text{test}}$, we return the prediction if and only if $\max{(\mathbf{\hat{y}}_{\text{test}})}\geq q$, otherwise we defer to the LM.
-In all experiments we set $\alpha=0.9$.
-If no satisfying $q$ can be found we defer on all test instances.
+During inference, when a probe predicts a behavior with probability vector $\mathbf{\hat{y}}_{\text{test}}$, we return the prediction if and only if $\max{(\mathbf{\hat{y}}_{\text{test}})}\geq q$, otherwise we defer to the LM.
+Unless specified otherwise, we set $\alpha=0.9$.
+If no satisfying $q$ can be found, we defer on all instances.
 We ablate $\alpha$ in Section 6.
-Using the conformal probes (Figure 2) significantly increases the performance across all datasets.
-For datasets where the total test consistency was already high (ARC with $88\%$ total test consistency), the conformal probes predict on large portions of the test set with consistency close to 90%.
-On datasets where total test consistency was poor (MedMCQA, MMLU, PiQA), the probes predict on fewer instances.
-When they do predict, they do so with higher consistencies than on the full test set.
-4 Accelerating Inference-Scaling with Conformal Probes
-Figure 3: Tradeoff between accuracy loss (CoT Accuracy - Accelerated CoT Accuracy) and inference cost when using our method to accelerate CoT Prompting in IID (Left) and OOD (Right) settings.
-In IID setting, Accuracy loss is negative on average, and never more than 1.34%, while inference cost reductions are 65% on average.
-In the OOD setting, inference cost reductions are higher at 81% on average, with a small accuracy loss of 2.3% on average.
-We extend the method described above to 18 other datasets spanning the tasks of Sentiment Analysis, Topic Analysis, Toxicity Detection and Fact Verification.
+Using conformal probes (Figure 2) significantly increases performance across all datasets.
+At $\alpha=0.8$, the probes have a test coverage of 40% and, for the covered datapoints, estimation consistency is 80.7%, showing that the conformal probes are well calibrated to the user-specified confidence level.
+Raising $\alpha$ to $0.9$ results in a reduced average coverage of 10%, but an increased average consistency of 89.2%.
+This allows users to flexibly trade off between coverage and consistency, enabling precise and trustworthy early warning systems that do not offer an opinion when confidence is low.
+Conformal Probe ($\alpha=0.9$) BERT
+Dataset Coverage Consistency Prec. Rec. Consistency Prec. Rec.
+SelfAware 100.0 98.0 94.5 94.5 93.3 91.7 84.4
+KnownUnkown 92.3 92.6 75.1 72.8 67.1 64.4 93.1
+WildJailbreak 100.0 94.0 90.0 91.0 83.4 73.8 33.3
+Table 1: Probes estimate whether the LM will fail to abstain when it should have with significantly higher precision and recall than a BERT baseline.
+On WildJailbreak, where 30% of malicious prompts are mistakenly complied with, using the probe as an early warning signal reduces successful jailbreaks to 2.7%.
+In comparison, using the BERT model would leave jailbreak success rate at 20%.
+4 Creating Early Warning Systems for Other Behaviors of Interest
+In this section, we demonstrate that probes are not limited to preemptively identifying degenerate behavior, such as failing to follow instructions.
+We show that conformal probes can serve as early warning systems for instances that will trigger failures to follow safety alignment, and those that will result in the model giving a low-confidence response.
+We describe the key experiment designs below, with exact prompts and details in the Appendix C.
+4.1 Detecting failures to follow safety alignment
+The deployment of LM-powered chatbots raises concerns of harmful behavior that is unaligned with human interests [28].
+One common setting where such behaviors arise is when the user requests information that does not exist, with models often hallucinating an answer regardless, misleading the user in the process [9].
+Another common occurrence is when the user themselves requests the model to help them complete malicious tasks, such as building a weapon of mass destruction [90].
+We collect two datasets (SelfAware [96], KnownUnknown [5]), which have answerable and unanswerable questions.
+The LM reasons as to whether the question is answerable, providing a response if it is and abstaining otherwise.
+Similarly, we collect malicious and benign prompts from WildJailbreak [45], and prompt the LM to refuse to comply with a request if it is malicious.
+Inspired by work in AI safety and alignment, we adopt the viewpoint that answering an unanswerable question or complying with a malicious request is disproportionately problematic [11].
+We filter our training set to only include the instances where the LM answers or complies with the request, and train probes to identify the cases where the LM should have abstained (i.e.
+fails to avoid answering an unanswerable question or falls victim to jailbreaking).
+During inference, if the probe is confident ($\alpha=0.9$) that the LM has failed to abstain when it should have, we override the LM and abstain instead.
+Results show (Table 1) that probes greatly enhance the safety of the LMs, outperforming the BERT baseline on consistency, precision and recall.
+Across all datasets, probes identify over 72% of the failures to abstain (over 90% for 2/3 datasets).
+This safety comes at little cost; when the probes force an abstention, they are correct over 86% of the time (on average).
+They make confident estimations on at least 92% of the samples, and maintain conformal estimation consistency of at least 92%.
+Concretely, on WildJailbreak the use of conformal probes reduces the percentage of successful jailbreaks from 30% to 2.7%, a significant safety improvement.
+4.2 Quantifying LM uncertainty
+Figure 3: Estimation consistency when preemptively predicting an LM’s confidence in its response.
+Numbers inside the bars in bold italics are the coverage for conformal probes that engage in selective prediction.
+Probes outperform all other methods, with the confidence estimation tasks proving more challenging for BERT.
+Conformal probes have lower coverage rates than the format following (Section 3) and safety alignment (Section 4.1) tasks.
+However, they maintain consistency at the user-defined confidence level, ensuring that their early warning signals are reliable.
+Language models are prone to hallucination [98] and often have gaps in their knowledge [38], making methods that estimate the trustworthiness of specific outputs vital.
+Recent work [48, 46] approaches this problem by using the model itself to quantify how certain it is in its answer, either by explicitly prompting it to verbalize its confidence or using the per-token perplexity of the output.
+If we detect an untrustworthy response, we may abstain from answering or escalate the query to a more capable model.
+An early warning system that preemptively detects such responses before they are even generated would allow us to take these actions more efficiently.
+We train conformal probes to preemptively estimate how confident the LM will be in its response using both of these uncertainty quantification methods on NaturalQA, MSMarco and TriviaQA.
+In the case of the perplexity measure, we consider the bottom $25\%$ of scores to be ‘high confidence,’ and the top $25\%$ to be ‘low confidence’ (discarding the rest).
+In both cases, the probes attempt to identify whether the model will be confident in its output.
+This task proves more challenging for all methods, however, probes continue to outperform both the random baseline and the much larger BERT model.
+Despite this, the conformal probes maintain high consistency, making the early warning and redirection system feasible in practice.
+5 Selective Inference-Scaling with Conformal Probes
+In this section, we show how our method can be used to accelerate systems that use CoT prompting with an LM to perform text classification.
+We collect 27 text classification datasets, spanning the tasks of Multiple Choice Question Answering, Sentiment Analysis, Topic Analysis, Toxicity Detection and Fact Verification.
 Details on dataset setup are in Appendix C.
-For all of these, we use Llama3.1-8B under CoT prompting (outputting an explanation before the final answer).
-We train a linear probe that uses the internal representation of the final input token at the 18th layer to predict the class that the CoT model will eventually output, and then perform conformal calibration.
-During inference, if the conformal probe is confident we use the probe estimation as the final answer.
+We use Llama3.1-8B under CoT prompting to perform text classification, i.e., outputting an explanation before the final class prediction.
+We train a linear probe that uses the internal representation of the final input token at the 18th layer (based on results from previous experiments) to predict the class that the CoT model will eventually output, and then perform conformal calibration.
+During inference, if the conformal probe is confident, we interrupt generation and use the probe estimation as the final answer.
 If not, we allow the model to continue its CoT generation and provide the final answer.
-We compare this to the vanilla CoT on two metrics—Accuracy Loss (Accuracy of CoT - Accuracy of Method) and Inference Cost Reduction (#CoT Forward Passes - #Method Forward Passes# CoT Forward $\frac{\#\text{CoT Forward Passes - \#Method Forward Passes}}{\text{\# CoT Forward Passes}}$).
-The results (Figure 3) show that the method is highly effective at reducing the inference cost with minimal cost to accuracy.
+We note that this method of acceleration is synergistic with other inference optimization algorithms such as Speculative Decoding [55] or architectural changes [27].
+This is because in the cases where the probe does not interrupt the LM, it does not influence the remaining forward passes, hence conformal probes can be flexibly added on top of any inference paradigm for further efficiency.
+We compare to vanilla CoT on two metrics—Accuracy Loss (Accuracy of CoT - Accuracy of Method) and Inference Cost Reduction (#CoT Forward Passes - $\frac{\#\text{CoT Forward Passes - \#Method Forward Passes}}$$\text${\# CoT Forward Passes}}).
+Figure 4: Tradeoff between accuracy loss and inference cost when using probes to accelerate CoT Prompting in IID and OOD settings.
+Using IID probes leads to a 1.34% accuracy loss (worst case) and a negative accuracy loss on average (i.e.
+accuracy improves).
+Inference cost reductions are 65% on average.
+Probes generalize to OOD data, inference cost reductions are higher at 81% on average, with a small accuracy loss of 2.3% on average.
+The results (Figure 4, in blue) show that the method is highly effective at reducing the inference cost with minimal cost to accuracy.
 The minimum inference cost reduction is $4.7\%$, with an average reduction of $65\%$ across all datasets.
@@
 For each test dataset of the MCQ and Sentiment tasks, we train and calibrate using data from every other dataset.
-The results (Figure 3) show that the probes do exhibit OOD generalization, suggesting the method may be applicable even when there are slight shifts between training and test distributions.
-5 Creating Early Warning Systems for Behaviors of Interest
-(a) Unanswerable Question Abstention
-(b) Bullet Point Format Following
-(c) JSON Format Following
-(d) Perplexity Confidence Estimation
-(e) Verbalized Confidence Estimation
-Figure 4: Probes can provide precise early warning signals for a range of important behaviours
-The phenomenon of probes containing information on the eventual output sequence of the LM is not limited to predicting text classification answers but applies to a wide range of behaviors that can be represented with discrete labels.
-In this section, we showcase several other use cases of this finding.
-We describe the key experiment designs below, with exact prompts in the Appendix C.
-LM Abstention: We collect two datasets (SelfAware [68], KnownUnknown [4]) which have both answerable and unanswerable questions.
-The LM is guided, via few-shot examples, to first reason out loud with respect to the question content and then either answer it or refuse to answer if the question seems to be unanswerable.
-Separately, we collect malicious and benign prompts from WildJailbreak [30], and prompt the LM to refuse to comply with a request if it is malicious.
-In both cases, the conformal probes try to estimate whether the LM will give an answer or refuse.
-Results show that probes confidently estimate (Figure 4(a)) the abstention decision on over 15% of KnownUnknown, 44% of SelfAware and 83% of WildJailbreak, with over 90% across datasets.
-Format Following Errors: We collect 3 QA datasets (NaturalQA [38], MSMarco [46], TriviaQA [32]) and prompt the LM to an answer in a specific format.
-One format requires the answer to be organized in exactly 3 bullet points, and the other requires a JSON output which contains pre-specified fields.
-The probes estimate whether the output will adhere to the format.
-Predicting format following errors in a JSON format proves more challenging than the bullet point format( 4(b)).
-However, even on the JSON task, the probes maintain a minimum consistency of 84% and have confidence on over 8% of instances on 2/3 datasets while deferring completely on the other.
-Low Confidence Output: On the same 3 QA datasets, we try to identify when the LM is less confident in its response.
-Following recent work [33, 31], we consider one setting where we explicitly prompt the model to verbalize its confidence, and another where we use the per-token perplexity of the output as a proxy for the confidence.
-In the case of the perplexity measure, we consider the bottom $25\%$ of scores to be ‘high confidence’, and the top $25\%$ to be ‘low confidence’ (discarding the rest).
-In both cases, the probes attempt to identify whether the model will be confident in its output.
-This is the most challenging task, probes typically make predictions on around 4% of test instances. However, consistency remains high, making the early warning system feasible in practice.
-The conformal probe-based early warning systems vary in their confidence across tasks and datasets.
-However, on a wide range of LM behaviors, they predict on sizable portions of the test set with consistently high consistency.
+The results (Figure 4, in orange) show that the probes do exhibit OOD generalization, suggesting the method may be applicable even when there are slight shifts between training and test distributions.
 6 Analysis and Ablations
-Figure 5: Left: Density plot of the correlation coefficient between the CoT model accuracy and the consistency of the probe estimate.
-The coefficient is often positive, indicating that instances where the probe inaccurately estimates CoT model are also instances where the model is more likely to be incorrect.
-Right: On CosmoQA, probes have near-random consistency with initial tokens.
-There is a sudden performance boost towards the final token.
-Confidence remains steady, suggesting poor calibration in early token probes.
 Why does accuracy improve when using probes?
-Seeking to explain the surprising increase in accuracy when accelerating CoT models (Figure 3), we plot (Figure 5) the correlation between the CoT models accuracy and the probes estimation consistency.
-It is generally a positive correlation, suggesting that on the instances where the probe estimates incorrectly, the CoT model is also more likely to output an incorrect answer.
-This minimizes the harm of an inconsistent estimation as sometimes making an inaccurate estimate is actively beneficial for overall accuracy.
+Seeking to explain the surprising increase in accuracy when accelerating CoT models (Figure 4), we plot (Figure 5, left) the correlation between the CoT model’s accuracy and the probe’s estimation consistency.
+It is generally positive, which suggests that incorrect probe estimation correlates with incorrect CoT generation.
+This minimizes the harm of an inconsistent estimation of the CoT decision, as the cases where estimates are likely to be inconsistent coincide with cases where a consistent estimate does not benefit accuracy.
+Figure 5: Left: Density plot of the correlation coefficient between the CoT LM accuracy and the consistency of the probe.
+The coefficient is often positive, indicating that instances where the probe inaccurately estimates the LM decision are also instances where the LM is more likely to be incorrect.
+Right: On CosmoQA, probes that use the initial tokens of the input prompt have near-random consistency, with performance increasing considerably as we use later tokens in the prompt.
 How early in the computation does the model show signs of the final behavior?
 We have shown that the final input token often contains sufficient information to predict output behavior, but at which token does this information start becoming clear and accessible?
-To investigate this, we collect the activations of the mid-layer from every input token (after the FewShot example tokens) and train a linear classifier to map any of these internal activations to the desired output behavior.
-Probing (Figure 5) using the embeddings of tokens in the first quarter of the question tokens leads to near-random performance.
-While consistency increases steadily as we use later tokens, a considerable jump is noticed when shifting to the final tokens.
-Surprisingly, confidence declines marginally as we use later tokens for probing.
-This suggests that using the earliest tokens for probing could result in inconsistent and poorly calibrated models, making it a poor design decision for this method.
-How does output length affect estimation accuracy?
-We measure the correlation between the estimation consistency of the probe and the token count of both the input (-0.41%) and output (-6.1%).
-While input length seems to have little effect, a higher output length is correlated with worse accuracy.
+To investigate this, we collect activations from every input token (after the FewShot example tokens) of the CSQA dataset and train a linear classifier to map these internal activations to the eventual CoT prediction.
+Probing (Figure 5, right) using the embeddings of tokens in the first quarter of the question leads to near-random performance, showing that this information is not readily accessible at initial tokens.
+Consistency increases steadily as we use later tokens, with a sharp rise around the final tokens.
+Figure 6: Left: Ablations on MMLU.
+Within 500 training datapoints, probes ($\alpha=0.9$) often approach the consistency and coverage they will attain when trained on the entire dataset.
+Layer ablations ($\alpha=0.9$) show that early layers perform poorly.
+Increasing the confidence threshold $\alpha$ has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
+Right: Estimation consistency increases with LM size, suggesting the method may scale favorably
+Ablations on probe parameters: We ablate the parameters involved in the design of the conformal probing system: number of training datapoints, probing layer and the conformal confidence threshold $\alpha$.
+We show results (Figure 6, left) for the MMLU dataset, however, all trends discussed hold for the majority of the datasets we explored, with more examples in Appendix B.
+We ablate the amount of data used to train the probe and find (Figure 6, top left) that probes can approach the consistency and coverage they will achieve over the entire dataset with only 500 datapoints, suggesting that the method is particularly data efficient.
+We explore (Figure 6, middle left) whether the specific layer being probed affects the performance of the probe.
+Our most general finding is that early layers offer poor estimation consistency.
+However, whether mid or late layers perform better is task and dataset-specific, and cannot be described generally.
+For more details, see Appendix B.
+Increasing the confidence required from the probes has predictable effects (Figure 6, bottom left).
+Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
+How does model scale affect performance?
+We varied the size of the LM used (Llama3 3B, 8B and 70B) and measured probe consistency across several layers.
+Encouragingly (Figure 6, right), the performance of the larger models is consistently better, suggesting that the information encoded in the internal activations is easier for the probe to ‘read’ when the model is more powerful.
+This bodes well for the methods’ ability to scale with models of increasing size and capability.
+What are some limitations of the probes?
+Other experiments we conducted suggest that probes are limited in the behaviors they can detect.
+Using the MCQ task, we tried to estimate whether the LM would output an incorrect answer.
+The probe accuracies were consistently near random.
+We hypothesize that since probes are simple linear classifiers, they can only detect patterns and use ‘knowledge’ that is well encoded in the LM activations.
+This suggests that probes struggle with output properties that cannot be identified by the LM itself (without external knowledge).
+We also observed that the correlation between the estimation consistency of the probe and the token count of the output is negative (-6.1%).
 This suggests probes struggle more with inputs that evoke longer outputs.
-Figure 6: Probes achieve high accuracy with as few as 1000 datapoints in most cases
-How much data does a probe need?
-We ablate the amount of data used to train the probe.
-We find (Figure 6) that often, fewer than 1000 datapoints is sufficient for the probe to achieve high estimation accuracy, suggesting that the method is particularly data efficient.
-Figure 7: Layer ablations show that early layers perform poorly, while the decision between mid and late layers is dataset specific
-Which layer should we use to probe?
-We explore (Figure 7) the specific layer being probed affected the performance of the probe.
-The most general finding is that the early layers offer poor estimation consistency.
-However, whether mid or late layers perform better is task and dataset-specific, and cannot be described generally.
-For more details see Appendix B.
-Figure 8: Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero
-How does varying $\alpha$ change performance?
-Increasing the confidence required from the probes has predictable effects (Figure 8).
-Increasing the confidence threshold has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
-Figure 9: Probe accuracy increases with model size, suggesting the method may scale favorably
-How does model scale affect performance?
-We varied the size of the LM used (Llama3 3B, 8B and 70B) and measure probe consistency across several layers.
-Encouragingly (Figure 9), the performance of the larger models are consistently better, suggesting that the information encoded in the internal activations is easier for the probe to ‘read’ when the model is more powerful.
-This bodes well for the methods’ ability to scale with models of increasing size and capability.
-What are the limits of the probes?
-Using the MCQ task, we tried to estimate whether the LM would output an incorrect answer.
-The probe accuracies were consistently poor and near random.
-This shows that, intuitively, the probes have limitations in the behaviors they can estimate.
-We hypothesize that since probes are simple linear classifiers, they can only detect patterns and use ‘knowledge’ that is well encoded in the internal activations being probed.
-This suggests that probes struggle with output properties that cannot be identified by the LM itself (without external knowledge).
 7 Related Works
 This work takes inspiration from recent methods that probe the hidden states of LMs to observe interpretable patterns [4, 51, 71, 36, 3] identify false statements [10, 56, 57, 97] and hallucinations [20, 84, 44].
-Our work also has connections to literature on early exiting during the forward pass of NN models, with works often using signals from the hidden states to prematurely exit with a prediction on the next token [65, 70, 66, 54, 27].
-Notably,  Schuster et al.
-[54] also uses a conformal prediction framework to give a provable error bound on the next token approximation.
-We are the first to show that the internal states can predict a range of behaviors, uninferable from the next token alone, before any of the output tokens are generated.
-Additionally, we are the first to show that conformal prediction can be used to create early warning systems for a wide range of behaviors like question abstention to format following errors.
-Our work advances research on understanding the nature of the information contained in the hidden states of LMs [50, 5, 47, 41, 61].
+These works use the internal states of every token, even the outputs, to produce signals for these phenomena.
+In contrast, our work shows that internal states can predict behaviors  before any output tokens are generated, suggesting that input token embeddings contain rich information on future LM behavior.
+Furthermore, while conformal prediction has been used with LMs to guarantee the quality of generated text [73], robotic trajectories [89, 74] and text classification [52, 17], we are the first to use it in conjunction with hidden state probes to create precise early warning systems.
+Our work also has connections to the literature on early exiting during the forward pass of NN models, with works often using signals from the hidden states to prematurely exit with a prediction on the future outputs [93, 99, 94, 80, 42, 70].
+While these works focus on the tokens predicted, we show a more general result: that internal states can predict the future behavior of LMs.
+These include behaviors such as whether or not the LM will output a high perplexity answer, or mistakenly comply with a malicious request, which cannot be inferred just by observing the output token sequence.
+While recent work has begun exploring whether LMs exhibit ‘introspective’ properties in the black box setting [13], or whether LMs can predict ‘global attributes’ of their response [72, 25], we are the first to show that when internal probes are deployed under conformal prediction they can be used to create early warning systems for a wide range of behaviors like question abstention to format following errors.
+Our work advances research on understanding the nature of the information contained in the hidden states of LMs [71, 6, 66, 58, 87, 62].
 Specifically, we show that the information contained in the hidden states is relevant not just to the next token, but to behaviors that manifest several tokens later during the LMs generation.
 8 Conclusion
-We show that a Language Model’s internal representation of input tokens alone contains vital information on the behavior of the LM over the entire output sequence.
-These signals can be used to create probes that serve as precise early warning or exit systems for LMs on a wide range of behaviors.
+We show that a language model’s hidden representation of input tokens alone contains vital information on the behavior of the LM over the entire output sequence.
+We train linear probes to read this information and use it in a conformal prediction framework to create precise early warning or exit systems for a wide range of LM behaviors, including degenerate behaviors, safety alignment failures and more.
+The conformal probes can preemptively identify, before a single token is generated, instances where an LM will fail to abstain from answering an unanswerable question, fall victim to jailbreaking, fail to follow output format specifications or give low-confidence responses.
 On 27 text classification datasets across 5 different tasks, the method can accelerate Chain-of-Thought prompting by 65% with little accuracy loss.
-The method can be used to create high-precision early warning systems that identify if an LM will refuse to answer a question, fall victim to jailbreaking, fail to follow output format specifications or give low-confidence responses.
-We show that the probes generalize to out-of-distribution test sets, and scale favorably to larger LMs.
+We show that the probes generalize to out-of-distribution test sets and scale favorably to larger LMs.
 Finally, we explore the limitations of the method, showing that the behavior of longer output sequences is harder to estimate and that tasks that require knowledge external to the model are particularly challenging.
-With the rising popularity of inference-time scaling methods, we hope our work can help ameliorate the growing computational cost of running LMs and provide more insight into the nature of the information contained in the hidden states of LMs.
+With the rising popularity of inference-time scaling methods, we hope our work can help ameliorate the growing computational cost of running LMs and provide more insight into the nature of the information contained in their hidden states.
+Acknowledgments
+We thank Robin Jia for his exceptional course on The Science of Large Language Models (accessible at https://robinjia.github.io/classes/fall2024-csci699.html) and his advice on the early stages of this project.
+We also acknowledge support from Open Philanthropy.
 Appendix A Confirming Robustness of Results
 A.1 Confirming Robustness to Choice of Model
-We conduct experiments on two other Language Models, to ensure that our results hold on Language Model families outside of the Llama3 series.
+We conduct experiments on two other language models to ensure that our results hold for language model families outside of the Llama3 series.
 We use Mistral-7B-Instruct-v0.3 from MistralAI [43] and DeepSeek-R1-Distill-Qwen-14B from DeepSeek [32].
@@
 A.2 Data Leakage
-Several datasets used in this study have been released before the creation of Llama3, and as such might have been pre-exposed to the model during training.
+Several datasets used in this study have been released before the creation of Llama3, and as such, might have been pre-exposed to the model during training.
 We detail our reasons as to why data leakage is not a concern for the conclusions that we draw from this study:
   • There are datasets in our study that have not been leaked.
-For example, the MMLU test set is used as an official test benchmark by LLama3[16] and WildJailbreak was released to the public only after the initial release of Llama3 models [30].
+For example, the MMLU test set is used as an official test benchmark by LLama3[26], and WildJailbreak was released to the public only after the initial release of Llama3 models [45].
 The results are strong on such datasets as well, showing that data leakage is not behind the performance of the probes.
   • Several tasks require the probes to identify behavior that fundamentally does not exist in text form on the internet.
-The bullets, JSON and Confidence tasks require the probes to determine how the model will behave, regardless of whether the model has seen the input text or not.
+The abstention failure, format following (bullets and JSON) and confidence estimation tasks require the probes to determine how the model will behave, regardless of whether the model has seen the input text or not.
 For example, whether or not the model was pretrained on MSMarco questions is immaterial for predicting whether or not it will output answers in a specific JSON format.
 Appendix B Extended Analysis
-B.1 Layer Ablation
+B.1 Training Datapoint Ablation
+Figure 9: Within 500 training datapoints, probes often approach the consistency and coverage ($\alpha=0.9$) they will attain when trained on the entire dataset.
+As a general trend (Figure 9), 500 training datapoints are usually sufficient for probes to achieve the maximum consistency and coverage that they will attain when trained on the entire dataset.
+B.2 Layer Ablation
 (a) Immediate saturation
@@
 A subset of the common patterns (Figure 10) shows that sometimes that apart from early layers being poor, it is unclear whether the middle or final layers are consistently better for accuracy or confidence.
+B.3 Confidence Threshold ablation
+Figure 11: Increasing the confidence threshold $\alpha$ has no effect until a point, after which conformal consistency increases while the coverage tends to zero.
+As a general trend (Figure 11), increasing the confidence threshold $\alpha$ has no impact until a point, after which it leads to a steady decline in coverage and a steady increase in estimation consistency.
 Appendix C Experiment Details
@@
 C.1.5 Toxicity Detection:
-We use two datasets provided by the same organization, JigsawToxicity  and a JigsawUnintendedBiasToxicity . Both datasets have scores for toxicity from a chatroom setting.
+We use two datasets provided by the same organization, JigsawToxicity and JigsawUnintendedBiasToxicity [1]. Both datasets have scores for toxicity from a chatroom setting.
 C.1.6 Others:
@@
 MCQ: For the MCQ tasks, we select two options by randomly sampling an incorrect answer from the provided options along with the correct option.
-This is because the MCQ datasets vary in terms of number of MCQ options, and for the out-of-distribution experiment we require all the datasets to have the same number of answer classes.
+This is because the MCQ datasets vary in terms of number of MCQ options, and for the out-of-distribution experiment we require all the datasets to have the same number of answer classes (note that the topic identification task below is multiclass, showing that our method is not limited to binary classification).
 The CoT model is asked to identify the correct answer option.
@@
 Hence, we instruct the model to comply with the request if it is not malicious, and abstain otherwise.
-In all cases, the probe is trained to detect whether or not the LM will abstain.
+We keep only the instances where the LM responds or complies, and train probes to detect whether or not the LM should have abstained (i.e.
+unanswerable question or malicious request).
+Format Following: Using an input prompt, we specify two different output formats a LM must obey.
+In both cases, the probe learns whether the LM will fail to follow format specifications or comply with them.
+  • Bullets: The output should be presented in 3 numbered bullet points, no fewer and no more
+  • JSON: The output should be presented in a JSON string with the following structure: {[‘short_answer’]: <str>, ‘entities’: List<str>, ‘references’: List<str>}
 Confidence Estimation: There are two settings for this task:
@@
   • Verbal Confidence: The LM is prompted to output an answer, along with a confidence score (either confident or unsure). The probe is trained to identify what the confidence score will be.
-Format Following: Using an input prompt, we specify two different output formats a LM must obey.
-In both cases, the probe learns whether the LM will fail to follow format specifications or comply with them.
-  • Bullets: The output should be presented in 3 numbered bullet points, no fewer and no more
-  • JSON: The output should be presented in a JSON string with the following structure: {[‘short_answer’]: <str>, ‘entities’: List<str>, ‘references’: List<str>}
-C.3 Prompts:
+C.3 Prompts and Inference:
 We have written separate prompts for each dataset to ensure the LM follows the instructions and outputs the text in a way that guarantees we can parse it and infer the behavior we seek to pre-emptively identify with the probes.
@@
  Answer:
-Each task and dataset has a different set of prompts, we have provided a link to the code https://github.com/DhananjayAshok/LMBehaviorEstimation. All prompts can be seen in the file data.py
+Each task and dataset has a different set of prompts. We have provided a link to the code, and all prompts can be seen in the file data.py
+All LM inference uses greedy decoding and is hence deterministic.
 C.4 Hardware:
@@
 The CPU on the cluster is an AMD EPYC 7502 32-Core Processor.
-Most experiments could be conducted with less than 16GB of RAM.
+Most experiments could be conducted with less than 16GB of GPU RAM.
+Generating the hidden states took around 2 hours for every 10,000 points, while training a hidden state probe takes fewer than 30 seconds.
 C.5 Replicability:
-To ensure that our code is easy to replicate and our method is easy to extend, we have provided open access to our code at https://github.com/DhananjayAshok/LMBehaviorEstimation
+To ensure that our code is easy to replicate and our method is easy to extend, we have provided open access to our code.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The main claims made in the abstract and introduction are that the internal states of the input tokens can predict the behavior of the LM on the entire output sequence as a whole, and that this can be used as a signal to create precise early warning systems for a variety of important behaviors.
+We show this through a set of experiments, where probes trained on the internal states successfully predict the future behavior of LMs, and can be used to preemptively detect instruction following errors, jailbreaking, etc.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: In the analysis section (Section 6), we dedicate a section to discuss the limitations of the probes.
+We detail how the kinds of behaviors the probes can detect have limits, and also show that certain kinds of inputs (ones that produce longer outputs) are harder for the probes to model.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14.
+Justification: While we use theoretical results from conformal prediction (see Section 2), these are not our theoretical results, and so we do not provide a proof.
+There is no original theoretical result in this work.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We have discussed the core implementation details in the Appendix C, including the dataset-specific set-up and hardware used.
+These details are a sufficient high-level guide to the experiments conducted in the paper, and allow a researcher to reproduce a similar version of the results even if they do not have access to our code.
+However, as detailed in the answer below, to ensure that the prompts, random seeds, generation parameters etc are transparent, we have released an anonymous version of our code base.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: We have provided an anonymous link to our code: https://anonymous.4open.science/r/LMBehaviorEstimation/.
+This code allows the reviewers (and future researchers) to fully reproduce our results.
+Since we use only open-sourced models with fixed random seeds for all stochastic pieces of code, we can guarantee a completely reproducible pipeline (given differences in hardware).
+Unfortunately, we cannot submit the data (specifically the hidden states used to train the probes) due to its size exceeding the supplemental material limits.
+However, the code will allow a complete reproduction of these hidden states, maintaining the reproducibility of our results.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We have provided a sufficient level of detail in the main text (Section 3) and the Appendix C to comprehend the results.
+While we have not provided full details as there are a large number of hyperparameters, the code base contains full details.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+[…]
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Appendix C explicitly details the hardware that was used to run the experiments, and gives an estimate on the compute/time required for each of them.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research does not involve human subjects and uses only publically accessible data.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49.
+Justification: A core motivation of this work is that of safety and early detection of harmful LM behavior.
+Through our experiments in Section 1, Section 4.1 and Section 4.2 we have shown that the deployment of our method could potentially have positive societal impacts by enabling more efficient and pre-emptive guardrails for harmful behaviors.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The code, data and probes that we released are all based on existing releases, and have little potential for misuse on their own.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All datasets and models used in this paper are done so in line with their stated licenses.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The new asset we introduce is a codebase, with documentation that enables reproduction and use.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification:
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification:
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79.
+Justification: This is a paper about LLMs, as such they are central to the research.
+We have outlined how we have used the LLMs clearly, and described all relevant design decisions.
+LLMs were not used in the writing of this manuscript, or in the creation of the core method/experiments in the paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.